# Photometric Inverse Rendering: Shading Cues Modeling and Surface Reflectance Regularization

### v1

@@ 1 Introduction
However, IRON [84] presents several shortcomings. Firstly, it assumes an ideal collocated camera-lighting arrangement, neglecting the complications posed by self-shadows that are frequently unfeasible in casual capture contexts, like smartphone use. Secondly, it fails to consider the diverse high-frequency inter-reflections characteristic of multi-view images captured with flashlight illumination. Such oversights can lead to inaccuracies in the estimation of diffuse albedo, as it allows self-shadows to distort the outcomes or unintentionally incorporates specular highlights, particularly within concave regions. Additionally, the absence of effective reflectance regularization in IRON undermines the precision of material decomposition.
In this work, we introduce a novel method that leverages rich shading information available in photometric images to achieve robust inverse rendering. Notably, our approach removes the necessity of co-locating the point light source with the camera, instead opting for a joint optimization of the light source’s position. This optimization accounts for the intricate interplay between the object’s geometry and the light source’s position, enabling our algorithm to effectively deduce the presence of self-shadows and significantly diminish their distorting impact on the resultant images. To accurately simulate the effect of inter-reflections, our method integrates an effective importance sampling strategy alongside a differentiable rendering layer. These techniques effectively reducing the unwanted blending of inter-reflections on material properties. To alleviate the ambiguity inherent from reflectance estimation, we incorporate a DINO [9] feature regularization into our inverse rendering framework. The self-supervised DINO method, by learning from extensive unlabeled datasets, captures image features encoding view-consistent contextual information across the scene, providing valuable information to understand the reflectance properties of different image regions and advancing the accuracy of the decomposition process.
In summary, our key contributions are as follows:
  • We propose a novel neural inverse rendering framework tailored for photometric images that jointly optimizes object shapes, materials and lighting, achieving accurate reflectance decomposition.
@@
  • We present a new dataset contains 5 scenes captured by a mobile phone in a darkroom. The number of images per scene ranges from 120 to 400.
@@ 2 Related Work
Many follow-up methods aim to improve the performance of NeRF on different surfaces types. Ref-NeRF [59] re-parameterizes NeRF’s outgoing radiance based on the reflection of the viewing vector with respect to the local normal vector, leading to improved rendering for specular surfaces [17, 90]. Some methods extend NeRF to handle more complex scenes containing mirror surfaces [57, 28, 77, 81, 20] and transparent objects [3, 48, 62].
@@
Efficiently computing inter-reflections, which involve tracing multiple bounces of rays, is a challenging problem in inverse rendering. Existing methods handle inter-reflections by assuming fixed illumination among multi-view images [53, 86, 14, 67, 74, 82, 73]. For example, InvRender [86] introduces an indirect illumination MLP to map a 3D point to its indirect incoming illumination, directly derived from the outgoing radiance field. However, our setup involves each image being illuminated under a different point light, making the existing approach unsuitable for our scenario.
To regularize the decomposition of reflectance, NeRFactor [85] learns a data prior for BRDFs by training an auto-encoder on the MERL dataset [43]. Some methods apply low-rank or vector-quantization regularization on the reflectance [88, 86, 89]. In comparison, we introduce a novel regularization without the need for additional training data by distilling the DINO [9] feature into the object’s surface. In the context of neural representation, DINO has been adopted in NeRF to scene editing  [27] and grouping semantic feature [26].
Figure 2: Method overview. Our method optimizes the light source position to account for self-shadows and models inter-reflection. The DINO features are injected into the networks of specular albedo and roughness to regularize the material decomposition.
@@
Different from the environment illumination, a point light model simplifies the image formation model and results in images with more high-frequency details, such as specular highlights, which significantly reduce ambiguity in inverse rendering [80, 69, 92, 58, 37]. Several methods have been proposed to improve the accuracy of inverse rendering by utilizing a point light model [5, 4, 71, 19, 8, 12, 30]. IRON [84] employs the NeuS method [63] to represent the surface and utilizes an edge-aware physics-based surface rendering for geometry refinement and materials estimation. However, IRON assumes an ideal collocated camera-lighting setup and overlooks self-shadows and inter-reflections. In contrast, our method addresses all these issues and explicitly regularizes on surface reflectance to achieve more accurate inverse rendering. While some methods compute shadows during optimization, they typically assume the environment illumination [85, 10, 60] or known point light positions [72, 36, 38, 56]. Differently, our method utilizes self-shadow cues to calibrate the light positions.
@@ 3.1 Overview
In this work, we focus on reconstructing object geometries, materials, and illumination conditions from multi-view images lit by a flashlight. Previous methods [84, 12] assume a collocated camera-light setup and optimize the scene with a simplified rendering model. Such an ideal setting is impractical to attain in our daily capture like mobile phone, we consider a more general fixed setup akin to a camera mounted on a mobile phone. Our method (Fig. 2) tackles the complexities introduced by self-shadows, inter-reflections, and ambiguities in reflectance estimation, which are prevalent issues in current techniques.
@@
where $f_{v}(\boldsymbol{w}_{i};\boldsymbol{x})$ indicates the visibility of light along $\boldsymbol{w}_{i}$ at $\boldsymbol{x}$ that models self-shadows in the rendered image, and $f_{ir}$ accounts for the residual effects attributed to inter-reflections.
@@ 3.2 Neural Scene Representation
where $T_{j}=\exp\left(-\sum_{q=1}^{j-1}\sigma_{q}\delta_{q}\right)$ denotes the accumulated transmittance at sampled point $\mathbf{r}(t_{j})$, and $\boldsymbol{c}_{j}$ represents the point’s color. We incorporate the unbiased density conversion method [42], translating SDF values into density representations for scene’s geometry.
@@ 3.3 Light Source Optimization
The flashlight is modeled as a white point light source, similar to the configurations in [84, 11]. We assume the photographic equipment is in rigid capture setup, such that the relative positions of the camera and point lights are fixed across images.
@@
where $\alpha_{j}$ is the discrete opacity values. We compute the visibility online so that the object geometry and light position can be jointly optimized.
@@ 3.4 Differentiable Inter-reflection Computation
Inter-reflection occurs when light reflects from one surface to another. We observe that the specular surfaces exhibit more pronounced inter-reflections, and the main source of indirect illumination for a surface point $\boldsymbol{x}$ comes from its reflective direction $\boldsymbol{w}_{r}$, which is the reflection of the view angle around the surface normal. To efficiently compute the indirect light, we sample multiple rays near $\boldsymbol{w}_{r}$ for indirect radiance calculation. Initially, we identify the secondary intersection point $\boldsymbol{x}^{\prime}$ where the indirect bounce meets the surface, and then we compute the incoming radiance $L_{\text{ind}}(\boldsymbol{w}_{i};\boldsymbol{x})$ as the outgoing radiance from $\boldsymbol{x}^{\prime}$. Radiance rendering of the secondary intersection point only takes into account the intense lighting from the flashlight and $\boldsymbol{x}^{\prime}$, excluding points occluded from the flashlight. The indirect illumination results from integrating all these incoming radiances over the upper-hemisphere around $\boldsymbol{x}$ surface normal.
To mitigate the artifacts of insufficient sampling, we blend the radiance from incoming direction $\boldsymbol{w}_{i}$ around the reflective view using a learnable scalar $\gamma$. One straightforward parameterization approach involves mapping the scalar from the point coordinates and the implicit geometric feature. However, we discovered that relying on point position and local geometry information often restricts the representation of varying inter-reflections, particularly in concave areas. A more effective strategy involves utilizing common physical properties (distance, view direction and roughness) to deduce dynamic indirect illumination. Consequently, we express the indirect illumination component $f_{ir}(\boldsymbol{x})$ as:
@@ 3.5 DINO Regularization
To reduce the inherent ambiguity of reflectance estimation, we introduce a novel reflectance regularization based on the distilled DINO feature field. DINO [9] displays inherent capabilities for object decomposition by training on diverse unlabeled data, and has been successfully distilled into 3D fields for radiance editing [27] and open-vocabulary object grouping [26]. Inspired by these methods, we propose to distill the DINO feature from 2D images to 3D surfaces (object geometry) to learn a fine composition and contextual information of the object, resulting in a more consistent decomposition of surface reflectance and materials. In our implementation, we distill the DINO feature to the initial geometry field by minimizing the loss function:
@@
where $\boldsymbol{p}$ denotes the pixel in the 2D images, $\boldsymbol{x}(\boldsymbol{p})$ indicates the corresponding 3D surface point derived by ray tracing. The distillation process is minimizing the square distance between the learnable DINO feature $\boldsymbol{f}_{\text{dino}}$ on surface and the ViTs pre-trained on 2D images. The distilled DINO features are incorporated into the networks of specular albedo and roughness, providing regularization to enhance the accuracy of material decomposition.
In our experiment, we found the resolution of the DINO feature will also influence the ability of distinguishing the composition of objects. Higher resolution DINO features can assist in achieving finer material decoupling. Empirically, we upsample the image by a scaling factor of two to extract DINO features with a higher resolution.
@@ 4.1 Datasets
The synthetic dataset comprises six objects. Four objects with a variety of shapes and materials, namely duck, maneki, horse, and dragon, are used in IRON [84]. To conduct thorough analysis, we adopt another two objects: marble bowl, a concave bowl with complex light effects, and armchair with self-shadows in multiple views.
We consider a practical non-collocated flashlight and camera setting, similar to a mobile phone setup, with the angle between the camera and flashlight to the object center set to about 3 degrees. We render 200 images from random views under a non-collocated flashlight via Mitsuba [22] for training. We also render 100 images together with their diffuse albedo maps, specular albedo maps and roughness maps for test images to evaluate the quality of novel view synthesis and material decomposition.
@@
Figure 4: Qualitative comparison of state-of-art methods and our method on the synthetic dataset. The materials of NeILF++[82] are Base Color, Metallic, Roughness defined by simplified Disney principled BRDF and others are using Mitsuba roughplastic model.
Table 2: Quantitative comparison on synthetic data. The predicted albedos are scaled to match the GT light intensity when evaluation.
@@ 4.2 Comparison with Existing Methods
Figure 6: Quantitative and visual Results of Light Optimization.
@@ 4.3 Results on Real Data
We compare our method with IRON on the challenging real dataset. Figure 5 showcases the rendered images and material decomposition results. Compared with IRON, our method exhibits fewer shadows and indirect illumination effects baked into the diffuse albedo, given the credit to our modeling of inter-reflection and lighting optimization.
@@ 4.4 Ablation Studies
To gain a deeper insight into the efficacy of our approach, we have conducted a thorough analysis of our method. We evaluate the inter-reflection modeling, lighting optimization, and DINO regularization, to validate our method.
@@
The synthetic data rendered for ablation study on inter-reflection is in collocated camera-lighting setting to overview self-shadows and different physical shader settings used in different methods. We ablate the inter-reflection calculation and compare the results in Table 3 and Fig. 7. Our method accurately estimates the materials and without the inter-reflection modeling, the predicted specular albedo has artifacts and diffuse albedo baked the indirect illumination.
Figure 9: Relighting results with materials estimated our method.
@@
Similarly, we evaluate the DINO regularization under the collocated camera-lighting setting for ablation study. Table 3 and Figure 8 (a) shows the quantitative and visual results respectively. Our decomposition results surpass other methods and we also show the DINO regularization is much better than some empirical regularization used in [86, 14]. The DINO feature regularization can ease the inherent difficulty of reflectance decomposition by grouping consistent contextual information across the scene.
@@ 4.5 Relighting Results
5 Conclusion and Discussion
In this paper, we present a new and effective inverse rendering approach for reconstructing object shapes, materials, and lighting from photometric images. Our method optimizes the light source position to account for self-shadows and employs an online strategy for modeling inter-reflections through a differentiable rendering layer. Additionally, we incorporate the DINO regularization to help the decomposition of surface reflectance. Extensive experiments on synthetic and real datasets demonstrate that our method can address misalignments between camera and light sources and surpass state-of-the-art methods in material decomposition. Due to its ability to finely compose objects, DINO regularization has the potential for application in other inverse rendering settings, such as those involving environment illumination.
@@ 5 Conclusion and Discussion
Our method has several constraints. Primarily, we overlooks the consistency of novel views captured by a moving flashlight in the geometry initialization stage. Secondly, the BRDF model we employ is tailored for solid reflective surfaces, thus our approach is not suitable for reflective surfaces. We leave these for our future works.

### v2

# PIR: Photometric Inverse Rendering with Shading Cues Modeling and Surface Reflectance Regularization

@@ 1 Introduction
However, IRON [84] presents several shortcomings. Firstly, it assumes an ideal collocated camera-lighting arrangement, neglecting the complications posed by self-shadows that are frequently unfeasible in casual capture contexts, like smartphone use. Secondly, it fails to consider the diverse high-frequency inter-reflections characteristic of multi-view images captured with flashlight illumination. Such oversights can lead to inaccuracies in the estimation of diffuse albedo, as it allows self-shadows to distort the outcomes or unintentionally incorporate specular highlights, particularly within concave regions. Additionally, the absence of effective reflectance regularization in IRON undermines the precision of material decomposition.
In this work, we introduce a novel method that leverages rich shading information available in photometric images to achieve robust inverse rendering. Notably, our approach removes the necessity of co-locating the point light source with the camera, instead opting for a joint optimization of the light source’s position. This optimization accounts for the intricate interplay between the object’s geometry and the light source’s position, enabling our algorithm to effectively deduce the presence of self-shadows and significantly diminish their distorting impact on the resultant images. To accurately simulate the effect of inter-reflections, our method integrates an effective importance sampling strategy alongside a differentiable rendering layer. These techniques effectively reducing the unwanted blending of inter-reflections on material properties. To alleviate the ambiguity inherent in reflectance estimation, we incorporate a DINO [9] feature regularization into our inverse rendering framework. The self-supervised DINO method, by learning from extensive unlabeled datasets, captures image features encoding view-consistent contextual information across the scene, providing valuable information to understand the reflectance properties of different image regions and advancing the accuracy of the decomposition process. Our key contributions are as follows:
  • We propose a novel neural inverse rendering framework tailored for photometric images that jointly optimizes object shapes, materials, and lighting, achieving accurate reflectance decomposition.
@@
  • We present a new dataset containing 5 scenes captured by a mobile phone in a darkroom. The number of images per scene ranges from 120 to 400.
@@ 2 Related Work
Many follow-up methods aim to improve the performance of NeRF on different surface types. Ref-NeRF [59] re-parameterizes NeRF’s outgoing radiance based on the reflection of the viewing vector with respect to the local normal vector, leading to improved rendering for specular surfaces [17, 90]. Some methods extend NeRF to handle more complex scenes containing mirror surfaces [57, 28, 77, 81, 20] and transparent objects [3, 48, 62].
@@
Efficiently computing inter-reflections, which involve tracing multiple bounces of rays, is a challenging problem in inverse rendering. Existing methods handle inter-reflections by assuming fixed illumination among multi-view images [53, 86, 14, 67, 74, 82, 73]. For example, InvRender [86] introduces a MLP to map a 3D point to its indirect incoming illumination, directly derived from the outgoing radiance field. However, our setup involves each image being illuminated under a different point light, making the existing approach unsuitable for our scenario.
To regularize the decomposition of reflectance, NeRFactor [85] learns the data prior for BRDFs by training an auto-encoder on the MERL dataset [43]. Some methods apply low-rank or vector-quantization regularization on the reflectance [88, 86, 89]. In comparison, we introduce a novel regularization without the need for additional training data by distilling the DINO [9] feature into the object’s surface. In the context of neural representation, DINO has been adopted in NeRF to scene editing  [27] and grouping semantic feature [26].
Figure 2: Method overview. Our method optimizes the light source position to account for self-shadows and model inter-reflection. The DINO features are injected into the networks of specular albedo and roughness to regularize the material decomposition.
@@
Different from the environment illumination, a point light model simplifies the image formation model and results in images with more high-frequency details, such as specular highlights, which significantly reduce ambiguity in inverse rendering [80, 69, 92, 58, 37]. Several methods have been proposed to improve the accuracy of inverse rendering by utilizing a point light model [5, 4, 71, 19, 8, 12, 30]. IRON [84] employs the NeuS method [63] to represent the surface and utilizes an edge-aware physics-based surface rendering for geometry refinement and materials estimation. However, IRON assumes an ideal collocated camera-lighting setup and overlooks self-shadows and inter-reflections. In contrast, our method addresses all these issues and explicitly regularizes surface reflectance to achieve more accurate inverse rendering. While some methods compute shadows during optimization, they typically assume the environment illumination [85, 10, 60] or known point light positions [72, 36, 38, 56]. Differently, our method utilizes self-shadow cues to calibrate the light positions.
@@ 3.1 Overview
In this work, we focus on reconstructing object geometries, materials, and illumination conditions from multi-view images lit by a flashlight. Previous methods [84, 12] assume a collocated camera-light setup and optimize the scene with a simplified rendering model. Such an ideal setting is impractical to attain in our daily capture, like a mobile phone; we consider a more general fixed setup akin to a camera mounted on a mobile phone. Our method (Fig. 2) tackles the complexities introduced by self-shadows, inter-reflections, and ambiguities in reflectance estimation, which are prevalent issues in current techniques.
@@
where $f_{v}(\boldsymbol{w}_{i};\boldsymbol{x})$ is the visibility of light along $\boldsymbol{w}_{i}$ at $\boldsymbol{x}$ that models self-shadows in the rendered image, and $f_{ir}$ accounts for the residual effects attributed to inter-reflections.
@@ 3.2 Neural Scene Representation
where $T_{j}=\exp\left(-\sum_{q=1}^{j-1}\sigma_{q}\delta_{q}\right)$ denotes the accumulated transmittance at sampled point $\mathbf{r}(t_{j})$, and $\boldsymbol{c}_{j}$ represents the point’s color. We incorporate the unbiased density conversion method [42], translating SDF values into density representations for the scene’s geometry.
@@ 3.3 Light Source Optimization
The flashlight is modeled as a white point light source, similar to the configurations in [84, 11]. We assume the photographic equipment is in a rigid capture setup, such that the relative positions of the camera and point lights are fixed across images.
@@
where $\alpha_{j}$ is the discrete opacity value. We compute the visibility online so that the object geometry and light position can be jointly optimized.
@@ 3.4 Differentiable Inter-reflection Computation
Inter-reflection occurs when light reflects from one surface to another. We observe that the specular surfaces exhibit more pronounced inter-reflections, and the main source of indirect illumination for a surface point $\boldsymbol{x}$ comes from its reflective direction $\boldsymbol{w}_{r}$, which is the reflection of the view angle around the surface normal. To efficiently compute the indirect light, we sample multiple rays near $\boldsymbol{w}_{r}$ for indirect radiance calculation. Initially, we identify the secondary intersection point $\boldsymbol{x}^{\prime}$ where the indirect bounce meets the surface, and then we compute the incoming radiance $L_{\text{ind}}(\boldsymbol{w}_{i};\boldsymbol{x})$ as the outgoing radiance from $\boldsymbol{x}^{\prime}$. Radiance rendering of the secondary intersection point only takes into account the intense lighting from the flashlight and $\boldsymbol{x}^{\prime}$, excluding points occluded from the flashlight. The indirect illumination results from integrating all these incoming radiances over the upper hemisphere around $\boldsymbol{x}$ surface normal.
To mitigate the artifacts of insufficient sampling, we blend the radiance from the incoming direction $\boldsymbol{w}_{i}$ around the reflective view using a learnable scalar $\gamma$. One straightforward parameterization approach involves mapping the scalar from the point coordinates and the implicit geometric feature. However, we discovered that relying on point position and local geometry information often restricts the representation of varying inter-reflections, particularly in concave areas. A more effective strategy involves utilizing common physical properties (distance, view direction, and roughness) to deduce dynamic indirect illumination. Consequently, we express the indirect illumination component $f_{ir}(\boldsymbol{x})$ as:
@@ 3.5 DINO Regularization
To reduce the inherent ambiguity of reflectance estimation, we introduce a novel reflectance regularization based on the distilled DINO feature field. DINO [9] displays inherent capabilities for object decomposition by training on diverse unlabeled data and has been successfully distilled into 3D fields for radiance editing [27] and open-vocabulary object grouping [26]. Inspired by these methods, we propose to distill the DINO feature from 2D images to 3D surfaces (object geometry) to learn a fine composition and contextual information of the object, resulting in a more consistent decomposition of surface reflectance and materials. In our implementation, we distill the DINO feature to the initial geometry field by minimizing the loss function:
@@
where $\boldsymbol{p}$ denotes the pixel in the 2D images, $\boldsymbol{x}(\boldsymbol{p})$ indicates the corresponding 3D surface point derived by ray tracing. The distillation process minimizes the square distance between the learnable DINO feature $\boldsymbol{f}_{\text{dino}}$ on the surface and the ViTs pre-trained on 2D images. The distilled DINO features are incorporated into the networks of specular albedo and roughness, providing regularization to enhance the accuracy of material decomposition.
In our experiment, we found that the resolution of the DINO feature also influences the ability to distinguish objects’ composition. Higher resolution DINO features can assist in achieving finer material decoupling. Empirically, we upsample the image by a scaling factor of two to extract DINO features with a higher resolution.
@@ 4.1 Datasets
The synthetic dataset comprises six objects. Four objects with a variety of shapes and materials, namely duck, maneki, horse, and dragon, are used in IRON [84]. To conduct a thorough analysis, we adopt another two objects: marble bowl, a concave bowl with complex light effects, and armchair with self-shadows in multiple views.
We consider a practical non-collocated flashlight and camera setting, similar to a mobile phone setup, with the angle between the camera and flashlight to the object center set to about 3 degrees. We render 200 images from random views under a non-collocated flashlight via Mitsuba [22] for training. We also render 100 images together with their diffuse albedo maps, specular albedo maps, and roughness maps for test images to evaluate the quality of novel view synthesis and material decomposition.
@@
Figure 4: Qualitative comparison of state-of-the-art methods and our method on the synthetic dataset. The materials of NeILF++[82] are Base Color, Metallic, Roughness defined by simplified Disney principled BRDF and others are using Mitsuba roughplastic model.
Table 2: Quantitative comparison on synthetic data. The predicted albedos are scaled to match the GT light intensity during evaluation.
@@ 4.2 Comparison with Existing Methods
Figure 6: Quantitative and Cisual Results of Light Optimization.
@@ 4.3 Results on Real Data
We compare our method with IRON on the challenging real dataset. Figure 5 showcases the rendered images and material decomposition results. Compared with IRON, our method exhibits fewer shadows and indirect illumination effects baked into the diffuse albedo, giving credit to our modeling of inter-reflection and lighting optimization.
@@ 4.4 Ablation Studies
To gain a deeper insight into the efficacy of our approach, we have conducted a thorough analysis of our method. We evaluate the inter-reflection modeling, lighting optimization, and DINO regularization to validate our method.
@@
The synthetic data rendered for the ablation study on inter-reflection is in a collocated camera-lighting setting to overview self-shadows and different physical shader settings used in different methods. We ablate the inter-reflection calculation and compare the results in Table 3 and Fig. 7. Our method accurately estimates the materials, and without the inter-reflection modeling, the predicted specular albedo has artifacts, and diffuse albedo baked the indirect illumination.
Figure 9: Relighting Results. Our method supports PBR relighting with the estimated materials.
@@
Similarly, we evaluate the DINO regularization under the collocated camera-lighting setting for the ablation study. Table 3 and Figure 8 (a) shows the quantitative and visual results respectively. Our decomposition results surpass other methods and we also show the DINO regularization is much better than some empirical regularization used in [86, 14]. The DINO feature regularization can ease the inherent difficulty of reflectance decomposition by grouping consistent, contextual information across the scene.
@@ 4.5 Relighting Results
5 Conclusion
In this paper, we present a new and effective inverse rendering approach for reconstructing object shapes, materials, and lighting from photometric images. Our method optimizes the light source position to account for self-shadows and employs an online strategy for modeling inter-reflections through a differentiable rendering layer. Additionally, we incorporate the DINO regularization to help the decomposition of surface reflectance. Extensive experiments on synthetic and real datasets demonstrate that our method can address misalignments between camera and light sources and surpass state-of-the-art methods in material decomposition. Thanks to its ability to precisely compose objects, DINO regularization holds promise for broader applications in inverse rendering, including scenarios involving environmental illumination.
@@ 5 Conclusion
Our method has several constraints. Primarily, we overlook the consistency of novel views captured by a moving flashlight in the geometry initialization stage. Secondly, the BRDF model we employ is tailored for solid reflective surfaces, thus our approach is not suitable for reflective surfaces. We leave these for our future work.
Acknowledgement This work was supported in part by NSFC with Grant No. 62293482, the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen-HK S&T Cooperation Zone. It was also partially supported by NSFC with Grant No. 62202409, the Shenzhen Science and Technology Program with Grant No. RCBS20221008093241052, the Shenzhen Outstanding Talents Training Fund 202002, the Guangdong Research Projects No. 2017ZT07X152 and No. 2019CX01X104, the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), and the Shenzhen Key Laboratory of Big Data and Artificial Intelligence (Grant No. ZDSYS201707251409055).